OpenStudy (anonymous):

Help with problem b-9+6b=30

OpenStudy (anonymous):

put like terms together

OpenStudy (anonymous):

b+6b=30+9

OpenStudy (anonymous):

then divide to get b by itself

OpenStudy (anonymous):

7b=39 b=39/7 b= 5 4/7